OpenStudy (anonymous):

The formula below gives A, the area of a circle with radius r. A(r)=πr^2 To the nearest hundredth, what is the value of A(8)? (Use 3.14 for π.) A)25.12 B)631.01 C)50.24 D)200.96

OpenStudy (anonymous):

\[A(r) = \pi (r^2)\] Here put r = 8 in this formula and solve further.. Can you do it??

OpenStudy (anonymous):

no this is hard :(

OpenStudy (anonymous):

See, \[A(8) = \pi \times 8^2\] Now can you solve?

OpenStudy (anonymous):

Tell me what is 8 squared??

OpenStudy (anonymous):

64

OpenStudy (anonymous):

Yes and what is the value of pi given in question?

OpenStudy (anonymous):

3.14

OpenStudy (anonymous):

YES ITS D :D

OpenStudy (anonymous):

so now, \[A(8) = (3.14) \times 64\] Can you go further to get the answer..

OpenStudy (anonymous):

ty ty tytty

OpenStudy (anonymous):

Yes you are right It is D..

OpenStudy (anonymous):

Welcome dear..